# “Fear will keep the local systems in line.”

### v2

@@ 7.1.4. Example
Indeed, the fibers are isomorphic to $\operatorname{Bun}_{G_{1}}$, where $G_{1}$ is a twisted form of $G^{\prime},G^{\prime}]$, the derived group of $G^{\prime}$, and the moduli stack of bundles for a simply-connected group is connected and simply-connected.
@@
Consider the stack $\operatorname{LS}_{M}$. It is quasi-smooth of virtual dimension
@@
The assertion of the lemma automatically extends to algebraic stacks. We apply it to the fibers of the map (7.4), i.e., to the stacks
@@
(7.8) $\Gamma(X,{\mathfrak{g}}_{{\mathcal{P}}_{G}}\otimes\omega_{X}).$
@@ 8.1. The notion of 2-Fourier-Mukai transform
(8.4) $\operatorname{2-FM}_{{\mathcal{Y}}_{1}\to{\mathcal{Y}}_{2}}:\operatorname{ShvCat}({\mathcal{Y}}_{1})\to\operatorname{ShvCat}({\mathcal{Y}}_{2}),\quad{\mathbf{C}}\mapsto(p_{2})_{*}((p_{1}^{*}({\mathbf{C}})_{{\mathcal{G}}_{1,2}}),$
@@
    see [Ga3, Sect. 3.1.3] (in loc. cit. it is denoted ${\mathbf{ind}}_{f}$);
@@
Note that the notion of being of 2-Fourier-Mukai type is in general asymmetric.
@@ 8.1.6. Example
Note also that the composition
$\operatorname{2-FM}_{\operatorname{pt}/\Gamma^{\vee}\to\operatorname{pt}/\Gamma}\circ\operatorname{2-FM}_{\operatorname{pt}/\Gamma\to\operatorname{pt}/\Gamma^{\vee}}$
is the identity endofunctor of $\operatorname{ShvCat}(\operatorname{pt}/\Gamma)$ up to the inversion involution on $\Gamma$.
8.2. 2-Fourier-Mukai transform and Poincaré duality
@@
Applying étale sheafification, we obtain a pairing
@@
(a) The pairing (8.10) is of 2-Fourier-Mukai type.
(b) The composition
$(\Gamma^{\vee}(1))^{\vee}(1)\simeq\Gamma.$
@@
Choose a point $x\in X$. Then we can split
$\operatorname{Ge}_{\Gamma}(X)\simeq\operatorname{Ge}_{\Gamma}(\operatorname{pt})\times\operatorname{pt}/H^{1}_{\operatorname{et}}(X,\Gamma)\times H^{2}_{\operatorname{et}}(X,\Gamma)$
@@
The pairing (8.10) splits as a product of:
@@
8.2.5.
@@
By the involutivity assertion in Theorem 8.2.3, we also obtain that for a $\Gamma$-gerbe ${\mathfrak{G}}_{\Gamma}$ on $X$, and the corresponding ${\mathbb{G}}_{m}$-gerbe ${\mathcal{G}}_{{\mathfrak{G}}_{\Gamma}}$ on $\operatorname{Ge}_{\Gamma^{\vee}(1)}(X)$, we have
(8.12) $\underline{{\mathbf{C}}}_{\Gamma}|_{{\mathfrak{G}}^{-1}_{\Gamma}}\simeq\mathbf{\Gamma}\left(\Gamma^{\vee}(1),(\underline{{\mathbf{C}}}_{\Gamma^{\vee}(1)})_{{\mathcal{G}}_{{\mathfrak{G}}_{\Gamma}}}\right).$
8.2.6.
@@
Unwinding the construction, we obtain:
@@
Unwinding the construction, we obtain:
@@ 8.3. Example: the usual Fourier-Mukai transform
8.3.1.
@@
${\operatorname{QCoh}}(\operatorname{Bun}_{T})\overset{\operatorname{FM}}{\simeq}{\operatorname{QCoh}}(\operatorname{Bun}_{T^{\vee}})\text{ and }{\operatorname{QCoh}}(\operatorname{Bun}_{T_{1}})\overset{\operatorname{FM}}{\simeq}{\operatorname{QCoh}}(\operatorname{Bun}_{T_{1}^{\vee}}).$
8.4. Some further compatibilities
In this subsection we review some further properties of the equivalence of Theorem 8.2.3 that will be needed for the proof of Theorem 8.5.8 below.
This subsection could be skipped on the first pass and returned to when necessary.
8.4.1.
Note that for any ${\mathfrak{G}}_{\Gamma^{\vee}(1)}\in\operatorname{Ge}_{\Gamma^{\vee}(1)}(X)$ we have the evaluation functor
8.4.2.
be the left adjoint of the evaluation functor
8.4.3.
Unwinding the constructions we obtain:
Lemma 8.4.4.
The following diagram commutes:
@@ 8.4. Some further compatibilities
8.4.5.
Let now
$0\to\Gamma_{1}\to\Gamma\to\Gamma_{2}\to 0$
be a short exact sequence of finite abelian groups, and let
$0\to\Gamma_{2}^{\vee}\to\Gamma^{\vee}\to\Gamma_{1}^{\vee}\to 0$
be the dual short exact sequence.
Generalizing (8.11), we have:
Lemma 8.4.6.
There is a canonical equivalence
@@ 8.5. Two sheaves associated with automorphic category
$\underline{\operatorname{D-mod}}^{Z_{G}}_{\frac{1}{2}}(\operatorname{Bun}_{G_{\operatorname{ad}}})\in\operatorname{ShvCat}(\operatorname{Ge}_{Z_{G}}(X))\text{ and }\underline{\operatorname{D-mod}}^{\pi_{1}({\check{G}})}_{\frac{1}{2}}(\operatorname{Bun}_{G})\in\operatorname{ShvCat}(\operatorname{Ge}_{\pi_{1}(G)}(X)).$
@@
$\operatorname{Bun}_{G,{\mathfrak{G}}_{Z_{G}}}:=\operatorname{Bun}_{G}\underset{\operatorname{Ge}_{Z_{G}}(X)}{\times}\{{\mathfrak{G}}_{Z_{G}}\}.$
@@
8.5.6.
We now consider $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G})$ as equipped with the spectral action of ${\operatorname{QCoh}}(\operatorname{LS}_{\check{G}})$. Since the stack $\operatorname{LS}_{\check{G}}$ is 1-affine, we can canonically attach to $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G})$ an object
@@
The short exact sequence of groups [^9] Note that ${\check{G}}_{\operatorname{sc}}$ is the Langlands dual of $G_{\operatorname{ad}}$.
@@
(8.22) ${\mathsf{p}}_{\pi_{1}({\check{G}})}:\operatorname{LS}_{\check{G}}\to\operatorname{Ge}_{\pi_{1}({\check{G}})}(X).$
@@
Combining Theorem 8.5.8 with the involutivity assertion in Theorem 8.2.3, we obtain:
@@
Indeed, the two theorems follow immediately from Lemma 8.2.8 and 8.2.10, respectively.
@@
The proposition will be proved in Sects. 8.7 and $8.$
8.6.8.
@@
8.7.1.
Let us return to the setting of Sect. 8.2. Let $\underline{x}$ be a point of $\operatorname{Ran}$}. Consider the spaces
@@
$\operatorname{Ge}_{\Gamma}(\mathcal{D}_{\underline{x}})_{\underline{x}}:=\operatorname{Fib}(\operatorname{Ge}_{\Gamma}(\mathcal{D}_{\underline{x}})\to\operatorname{Ge}_{\Gamma}(\mathcal{D}^{\times}_{\underline{x}}),$
@@
8.7.3.
@@
(8.34) $\operatorname{Ge}_{\Gamma}(\mathcal{D}_{\underline{x}})_{\underline{x}}\times\operatorname{Ge}_{\Gamma^{\vee}(1)}(\mathcal{D}_{\underline{x}})\to\operatorname{Ge}_{{\mathbb{G}}_{m}}(\operatorname{pt}),$
and it is easy to show that (8.34) is also of Fourier-Mukai type.
@@
8.7.4.
@@
The commutative group structure on the space of gerbes gives rise to a map
@@
There exists a symmetric monoidal functor
@@
8.7.10.
The constructions in Sect. 8.7.5 have immediate counterparts for the global Hecke stack
@@
Note that the space $\operatorname{Ge}_{\pi_{1}({\check{G}})}(X)_{x}$ is canonically the discrete set $\pi_{1}({\check{G}})(-1)$, which is in bijection with the set of characters of $Z_{G}$. Denote the element corresponding to our gerbe by $\chi$.
8.8.2.
We claim now that the proof of the spectral decomposition theorem in [Ga4, Sect. 11.1] applies in the current context. Recall that in loc. cit. the proof was based on considering the localization functor
@@
The key point in the proof of [Ga4, Theorem 4.5.2] is the fact (going back to [BD] and reviewed in [GLC2, Sects. 15-16]) that the functor
@@
Now, by the same principle as in [Ga4, Theorem 10.3.4] (see also [GLC2, Sects. 15-16]), we obtain that the functor $\operatorname{Loc}_{G_{\operatorname{ad}},\underline{x},\chi}$ factors as

### v3

@@ 7.1.4. Example
Indeed, the fibers are isomorphic to $\operatorname{Bun}_{G_{1}}$, where $G_{1}$ is a twisted form of $G^{\prime},G^{\prime}]$, the derived group of $G^{\prime}$, and the moduli stack of bundles for a simply-connected group is connected and simply-connected (see [BMP, Corollary 3.4]).
@@
Consider the stack $\operatorname{LS}_{M}$. It is quasi-smooth of virtual dimension [^16] See, e.g., [AG, Proposition 10.4.5].
@@
Corollary 7.3.11.
Let $\mathcal{Y}$} be a quasi-smooth algebraic stack of virtual dimension [^17] Recall that for a quasi-smooth algebraic stack $\mathcal{Y}$}, its dimension/virtual dimension are defined as follows: for a smooth cover $Y\to{\mathcal{Y}}$ with $Y$ a scheme, the dimension/virtual dimension of $\mathcal{Y}$} equals that of $Y$ minus the dimension of the fibers. $d$. Suppose that $m$ is an integer such that for all field-valued points $y\in{\mathcal{Y}}$ we have
$\dim(H^{-1}(T^{*}_{y}({\mathcal{Y}})))\leq m.$
Then $\dim({\mathcal{Y}})\leq d+m$.
We apply the corollary to the fibers of the map (7.4), i.e., to the stacks
@@
(7.8) $\Gamma(X,{\mathfrak{g}}_{{\mathcal{P}}_{G}}\otimes\omega_{X})$
(see [AG, Corollary 10.5.5] [^18] At the classical level, it is clear that connections on a given $G$-bundle on a curve form a torsor over the space of (Ad-twisted) $\check{{\mathfrak{g}}}$}-valued 1-forms. However, in derived algebraic geometry, this requires some care, hence the reference.).
@@ 8.1. The notion of 2-Fourier-Mukai transform
(8.4) $\operatorname{2-FM}_{{\mathcal{Y}}_{1}\to{\mathcal{Y}}_{2}}:\operatorname{ShvCat}({\mathcal{Y}}_{1})\to\operatorname{ShvCat}({\mathcal{Y}}_{2}),\quad\underline{{\mathbf{C}}}\mapsto(p_{2})_{*}((p_{1}^{*}(\underline{{\mathbf{C}}})_{{\mathcal{G}}_{1,2}}),$
@@
    see [Ga3, Sect. 3.1.3] (in loc. cit. it is denoted ${\mathbf{coind}}_{f}$);
@@
Note that the notion of being of 2-Fourier-Mukai type is a priori asymmetric.
@@ 8.1.6. Example
8.1.7.
Let us return to the general setting of Sect. 8.1.2. Let us make the following assumptions (essentially, on the geometric properties of ${\mathcal{Y}}_{1})$:
  • For $\underline{{\mathbf{D}}}\in\operatorname{ShvCat}({\mathcal{Y}}_{1}\times{\mathcal{Y}}_{2})$, the functor
    $(p_{2})^{*}\circ(p_{2})_{*}(\underline{{\mathbf{D}}})\to{\mathbf{D}}$
    admits a left adjoint;
  • For $\underline{{\mathbf{C}}}_{2}\in\operatorname{ShvCat}({\mathcal{Y}}_{2})$, the functor
    $\underline{{\mathbf{C}}}_{2}\to(p_{2})_{*}\circ(p_{2})^{*}(\underline{{\mathbf{C}}}_{2})$
    admits a left adjoint.
Note that this case, the above left adjoints provide a unit and counit, making $(p_{2})^{*}$ the right adjoint of $(p_{2})_{*}$ (i.e., the usual $((p_{2})^{*},(p_{2})_{*})$-adjunction is ambidexterous.
In particular, the new $((p_{2})_{*},(p_{2})^{*})$-adjunction gives rise to an adjunction [^19] In the formula below, $\operatorname{2-FM}_{{\mathcal{Y}}_{1}\to{\mathcal{Y}}_{2}}$ is defined using the initial gerbe ${\mathcal{G}}_{1,2}$, and $\operatorname{2-FM}_{{\mathcal{Y}}_{2}\to{\mathcal{Y}}_{1}}$ is defined using its inverse.
$(\operatorname{2-FM}_{{\mathcal{Y}}_{1}\to{\mathcal{Y}}_{2}},\operatorname{2-FM}_{{\mathcal{Y}}_{2}\to{\mathcal{Y}}_{1}}).$
In particular, we obtain:
Lemma 8.1.8.
Under the above circumstances, if $\operatorname{2-FM}_{{\mathcal{Y}}_{1}\to{\mathcal{Y}}_{2}}$ is an equivalence, then so is $\operatorname{2-FM}_{{\mathcal{Y}}_{2}\to{\mathcal{Y}}_{1}}$, and
$\operatorname{2-FM}_{{\mathcal{Y}}_{2}\to{\mathcal{Y}}_{1}}\circ\operatorname{2-FM}_{{\mathcal{Y}}_{1}\to{\mathcal{Y}}_{2}}\simeq\operatorname{Id}.$
8.2. Some compatibilities
8.2.1.
In this subsection, we will assume that ${\mathcal{Y}}_{1}:={\mathcal{H}}_{1}$ and ${\mathcal{Y}}_{2}:={\mathcal{H}}_{2}$ are commutative group-prestacks, and the map (8.3) is bilinear.
For a point $h_{1}\in{\mathcal{H}}_{1}$ (resp., $h_{2}\in{\mathcal{H}}_{2})$ let ${\mathcal{G}}_{h_{1}}$ (resp., ${\mathcal{G}}_{h_{2}}$) denote the corresponding ${\mathbb{G}}_{m}$-gerbe on ${\mathcal{H}}_{2}$
Let $\underline{{\mathbf{C}}}_{1}$ be an object of $\operatorname{ShvCat}({\mathcal{H}}_{1})$ and set $\underline{{\mathbf{C}}}_{2}:=\operatorname{2-FM}_{{\mathcal{H}}_{1}\to{\mathcal{H}}_{2}}({\mathbf{C}}_{2})$. Denote also
${\mathbf{C}}_{1}:=\mathbf{\Gamma}({\mathcal{H}}_{1},\underline{{\mathbf{C}}}_{1})\text{ and }{\mathbf{C}}_{2}:=\mathbf{\Gamma}({\mathcal{H}}_{2},\underline{{\mathbf{C}}}_{2}).$
8.2.2.
Note that for $h_{2}\in{\mathcal{H}}_{2}$ we have:
(8.10) $\underline{{\mathbf{C}}}_{2}|_{h_{2}}\simeq\mathbf{\Gamma}({\mathcal{H}}_{1},(\underline{{\mathbf{C}}}_{1})_{{\mathcal{G}}_{h_{2}}}),$
where:
  • $(-)|_{y_{2}}$ denotes the fiber of a given sheaf of categories at $y_{2}$;
  • $(-)_{{\mathcal{G}}}$ denotes the twist of a given sheaf of categories by a ${\mathbb{G}}_{m}$-gerbe.
As particular case of (8.10), we have:
(8.11) $\underline{{\mathbf{C}}}_{2}|_{{\mathbf{1}}_{{\mathcal{H}}_{2}}}\simeq{\mathbf{C}}_{1}.$
8.2.3.
Looping the pairing (8.3) along ${\mathcal{H}}_{2}$ we obtain a pairing
${\mathcal{H}}_{1}\times{\mathcal{A}}_{2}\to\operatorname{pt}/{\mathbb{G}}_{m}.,$
where
${\mathcal{A}}:=\operatorname{Aut}_{{\mathcal{H}}_{2}}({\mathbf{1}}_{{\mathcal{H}}_{2}}).$
In particular, a point $a\in{\mathcal{A}}$ gives rise to a line bundle, to be denoted ${\mathcal{L}}_{a}$ on ${\mathcal{H}}_{1}$.
Unwinding, we obtain:
Lemma 8.2.4.
Under the identification of (8.11), the action of $a$ on ${\mathbf{C}}_{2}|_{\mathbf{0}_{{\mathcal{H}}_{2}}}$ corresponds to the endofunctor of ${\mathbf{C}}_{1}$, given by tensoring wity ${\mathcal{L}}_{a}$.
8.2.5.
Consider now the group
$\Omega:=\operatorname{Aut}_{{\mathcal{A}}}({\mathbf{1}}_{\mathcal{A}}).$
Note that the group $\Omega$ acts on ${\mathcal{O}}_{{\mathcal{H}}_{1}}\in{\operatorname{QCoh}}({\mathcal{H}}_{1})$.
Let $\omega$ be an idempotent in the category ${\mathop{\operatorname{\rm Rep}}}(\Omega)$. Note if $\mathbf{d}$} is an object of a DG category, equipped with an action of $\Omega$, we can attach to it a direct summand ${\mathbf{d}}_{\omega}\subset{\mathbf{d}}$.
In particular, to $\omega$ there corresponds a connected component $({\mathcal{H}}_{1})_{\omega}$ of ${\mathcal{H}}_{1}$.
Unwinding, we obtain:
Lemma 8.2.6.
Under the identification of (8.11), the direct summand
${\mathbf{C}}_{1,\omega}:=\Gamma(({\mathcal{H}}_{1})_{\omega},\underline{{\mathbf{C}}}_{1})\subset\Gamma({\mathcal{H}}_{1},\underline{{\mathbf{C}}}_{1})={\mathbf{C}}_{1}$
corresponds to the direct summand
$(\underline{{\mathbf{C}}}_{2}|_{{\mathbf{1}}_{{\mathcal{H}}_{2}}})_{\omega}\subset\underline{{\mathbf{C}}}_{2}|_{{\mathbf{1}}_{{\mathcal{H}}_{2}}}.$
8.2.7.
Let us suppose that the assumptions from Sect. 8.1.7 hold. Unwinding, we obtain:
Lemma 8.2.8.
For $\underline{{\mathbf{C}}}_{1}\in\operatorname{ShvCat}({\mathcal{H}}_{1})$, the following diagram commutes:
[…]
where:
  • $\operatorname{ev}_{h_{2}}$ denotes the natural evaluation functor
    $\mathbf{\Gamma}({\mathcal{H}}_{2},\underline{{\mathbf{C}}}_{2})\to\underline{{\mathbf{C}}}_{2}|_{h_{2}}$
    for $\underline{{\mathbf{C}}}_{2}\in\operatorname{ShvCat}({\mathcal{H}}_{2})$;
  • $(\operatorname{ev}|_{{\mathbf{1}}_{{\mathcal{H}}_{1}}})^{L}$ is the left adjoint of the similarly defined evaluation functor on ${\mathcal{H}}_{1}$, where we note that
    $(\underline{{\mathbf{C}}}_{1})_{{\mathcal{G}}_{h_{2}}}|_{{\mathbf{1}}_{{\mathcal{H}}_{1}}}\simeq\underline{{\mathbf{C}}}_{1}|_{{\mathbf{1}}_{{\mathcal{H}}_{1}}},$
    since ${\mathcal{G}}_{h_{2}}$ is trivialized at ${\mathbf{1}}_{{\mathcal{H}}_{1}}$;
  • The lower left vertical arrow is the unit of the $(\operatorname{2-FM}_{{\mathcal{H}}_{1}\to{\mathcal{H}}_{2}},\operatorname{2-FM}_{{\mathcal{H}}_{2}\to{\mathcal{H}}_{1}})$-adjunction.
8.3. The 2-Fourier-Mukai transform and Verdier duality
@@
Composing with $\mu_{\infty}\to{\mathbb{G}}_{m}$ and applying étale sheafification, we obtain a pairing
@@
The pairing (8.12) is of 2-Fourier-Mukai type.
@@
Recall that we can think of $\operatorname{Ge}_{\Gamma}(X)$ as
$B^{2}(\operatorname{C}^{\cdot}(X,\Gamma))_{\operatorname{et}},$
and similarly for $\operatorname{Ge}_{\Gamma^{\vee}(1)}(X)$.
Choose a pair of points $x_{1},x_{2}\in X$; denote $U_{i}=X-x_{i}$. Restriction to $x_{1}$ gives rise to an isomorphism
$\operatorname{C}^{\cdot}(X,\Gamma)\simeq\operatorname{C}^{\cdot}(X;x_{1},\Gamma)\times\Gamma,$
where we note that
$\operatorname{C}^{\cdot}(X;x_{1},\Gamma)\simeq\operatorname{C}_{c}^{\cdot}(U_{1},\Gamma).$
The inclusion of $x_{2}$ gives rise to an isomorphism
$\operatorname{C}^{\cdot}(X;x_{1},\Gamma)\simeq\operatorname{C}^{\cdot}(U_{2};x_{1},\Gamma)\times\Omega^{2}(\Gamma(-1)),$
where $\Omega$ is the functor of loops on spectra, and where we note that
$\Omega^{2}(\Gamma(-1))\simeq\Omega^{2}(H^{2}_{\operatorname{et}}(X,\Gamma))\simeq\operatorname{C}^{\cdot}(X;U_{2},\Gamma)$
@@
$\operatorname{C}^{\cdot}(U_{2};x_{1},\Gamma)\simeq\Omega(H^{1}_{\operatorname{et}}(X,\Gamma)).$
Altogether we obtain an identification
$\operatorname{C}^{\cdot}(X,\Gamma)\simeq\Gamma\times\Omega(H^{1}_{\operatorname{et}}(X,\Gamma))\times\Omega^{2}(\Gamma(-1)),$
and hence
$\operatorname{Ge}_{\Gamma}(X)\simeq\operatorname{Ge}_{\Gamma}(\operatorname{pt})\times\operatorname{pt}/H^{1}_{\operatorname{et}}(X,\Gamma)\times H^{2}_{\operatorname{et}}(X,\Gamma).$
Similarly, we obtain
@@
Under this identification, the pairing (8.12) splits as a product of:
@@
8.3.5.
It is easy to see that the prestack $\operatorname{Ge}_{\Gamma}(X)$ satisfies the assumptions of Sect. 8.1.7. Hence, combining Theorem 8.3.3 and Lemma 8.1.8, we obtain:
Corollary 8.3.6.
The composition
$(\Gamma^{\vee}(1))^{\vee}(1)\simeq\Gamma.$
8.4. Some further compatibilities
In this subsection we will specialize the discussion in Sect. 8.2 to the special case of
${\mathcal{H}}_{1}:=\operatorname{Ge}_{\Gamma}(X)\text{ and }{\mathcal{H}}_{2}:=\operatorname{Ge}_{\Gamma^{\vee}(1)}(X).$
This is done primarily in order to have a convenient reference in subsequent subsections. The material here should be skipped and returned to when necessary.
8.4.1.
@@ 8.4. Some further compatibilities
8.4.2.
By Corollary 8.3.6, we also obtain that for a $\Gamma$-gerbe ${\mathfrak{G}}_{\Gamma}$ on $X$, and the corresponding ${\mathbb{G}}_{m}$-gerbe ${\mathcal{G}}_{{\mathfrak{G}}_{\Gamma}}$ on $\operatorname{Ge}_{\Gamma^{\vee}(1)}(X)$, we have
(8.14) $\underline{{\mathbf{C}}}_{\Gamma}|_{{\mathfrak{G}}^{-1}_{\Gamma}}\simeq\mathbf{\Gamma}\left(\operatorname{Ge}_{\Gamma^{\vee}(1)}(X),(\underline{{\mathbf{C}}}_{\Gamma^{\vee}(1)})_{{\mathcal{G}}_{{\mathfrak{G}}_{\Gamma}}}\right).$
8.4.3.
Let now
$0\to\Gamma_{1}\to\Gamma\to\Gamma_{2}\to 0$
be a short exact sequence of finite abelian groups, and let
$0\to\Gamma_{2}^{\vee}\to\Gamma^{\vee}\to\Gamma_{1}^{\vee}\to 0$
be the dual short exact sequence.
Generalizing (8.13) and (8.14), we have:
Lemma 8.4.4.
There is a canonical equivalence
8.4.5.
@@
The following is a particular case of (8.2.6):
@@
The followings is a particular case of Lemma 8.2.4:
@@
8.4.10.
For ${\mathfrak{G}}_{\Gamma^{\vee}(1)}\in\operatorname{Ge}_{\Gamma^{\vee}(1)}(X)$, denote by
$\operatorname{ev}|_{{\mathfrak{G}}_{\Gamma^{\vee}(1)}}:{\mathbf{C}}_{\Gamma^{\vee}(1)}\to(\underline{{\mathbf{C}}}_{\Gamma^{\vee}(1)})|_{{\mathfrak{G}}_{\Gamma^{\vee}(1)}}$
the canonical evaluation functor
8.4.11.
be the left adjoint of the evaluation functor
8.4.12.
As a particular case of Lemma 8.2.8, we obtain:
Lemma 8.4.13.
The following diagram commutes:
[…]
@@ 8.5. Example: the usual Fourier-Mukai transform
This subsection is included in order to illustrate how the 2-Fourier-Mukai transform works; this material is not needed in the rest of the paper.
8.5.1.
@@
${\operatorname{QCoh}}(\operatorname{Bun}_{T})\overset{\operatorname{FM}}{\simeq}{\operatorname{QCoh}}(\operatorname{Bun}_{T^{\vee}})\text{ and }{\operatorname{QCoh}}(\operatorname{Bun}_{T_{1}})\overset{\operatorname{FM}}{\simeq}{\operatorname{QCoh}}(\operatorname{Bun}_{T_{1}^{\vee}}),$
combined with the following observation:
The composition
@@
is canonically trivial. Furthermore, the resulting map
$\operatorname{Bun}_{T_{1}}\times\operatorname{Bun}_{T^{\vee}_{1}}\to\operatorname{Bun}_{T}\times\operatorname{Bun}_{T^{\vee}_{1}}\to\operatorname{pt}/{\mathbb{G}}_{m}$
equals the Weil pairing.
@@ 8.6. Two sheaves associated with automorphic category
$\underline{\operatorname{D-mod}}^{Z_{G}}_{\frac{1}{2}}(\operatorname{Bun}_{G_{\operatorname{ad}}})\in\operatorname{ShvCat}(\operatorname{Ge}_{Z_{G}}(X))\text{ and }\underline{\operatorname{D-mod}}^{\pi_{1}({\check{G}})}_{\frac{1}{2}}(\operatorname{Bun}_{G})\in\operatorname{ShvCat}(\operatorname{Ge}_{\pi_{1}({\check{G}})}(X)).$
@@
$\operatorname{Bun}_{G,{\mathfrak{G}}_{Z_{G}}}:=\operatorname{Bun}_{G_{\operatorname{ad}}}\underset{\operatorname{Ge}_{Z_{G}}(X)}{\times}\{{\mathfrak{G}}_{Z_{G}}\}.$
@@
8.6.6.
We now consider $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G})$ as equipped with the spectral action of ${\operatorname{QCoh}}(\operatorname{LS}_{\check{G}})$. Since the stack $\operatorname{LS}_{\check{G}}$ is 1-affine (say, by [Ga3, Theorem 2.2.4]), we can canonically attach to $\operatorname{D-mod}_{\frac{1}{2}}(\operatorname{Bun}_{G})$ an object
@@
The short exact sequence of groups
@@
(8.24) ${\mathsf{p}}_{\pi_{1}({\check{G}})}:\operatorname{LS}_{\check{G}}\to\operatorname{Ge}_{\pi_{1}({\check{G}})}(X),$
where we note that ${\check{G}}_{\operatorname{sc}}$ is the Langlands dual of $G_{\operatorname{ad}}$.
@@
Combining Theorem 8.6.8 with Corollary 8.3.6, we obtain:
@@
Indeed, the two theorems follow immediately from Lemma 8.4.7 and 8.4.9, respectively.
@@
The proposition will be proved in Sects. 8.8 and 8.9.
8.7.8.
@@
8.8.1.
Let us return to the setting of Sect. 8.3. Let $\underline{x}$ be a point of $\operatorname{Ran}$}. Consider the spaces
@@
$\operatorname{Ge}_{\Gamma}(\mathcal{D}_{\underline{x}})_{\underline{x}}:=\operatorname{Fib}(\operatorname{Ge}_{\Gamma}(\mathcal{D}_{\underline{x}})\to\operatorname{Ge}_{\Gamma}(\mathcal{D}^{\times}_{\underline{x}})),$
@@
In particular, we obtain that for $\underline{x}=x$ being a singleton, we have:
(8.36) $\operatorname{Ge}_{\Gamma}(\mathcal{D}_{\underline{x}})_{\underline{x}}\simeq\Gamma(-1),$
where the right-hand side is viewed as an algebro-geometrically discrete finite set.
8.8.3.
@@
(8.37) $\operatorname{Ge}_{\Gamma}(\mathcal{D}_{\underline{x}})_{\underline{x}}\times\operatorname{Ge}_{\Gamma^{\vee}(1)}(\mathcal{D}_{\underline{x}})\to\operatorname{Ge}_{{\mathbb{G}}_{m}}(\operatorname{pt}).$
@@
Remark 8.8.4.
Although we do not need it in what follows, we remark that the pairing (8.37) is also of Fourier-Mukai type. When $\underline{x}=x$ is a singleton, this follows immediately from the identifications
$\operatorname{Ge}_{\Gamma}(\mathcal{D}_{\underline{x}})_{\underline{x}}\simeq\Gamma(-1)\text{ and }\operatorname{Ge}_{\Gamma^{\vee}(1)}(\mathcal{D}_{\underline{x}})\simeq\Gamma^{\vee}(1).$
8.8.5.
@@
The commutative group structure on the space of gerbes gives rise to a map [^21] For a map of commutative group-objects $\mathcal{A}$$\to{\mathcal{B}}$, there is a canonical isomorphism ${\mathcal{A}}\underset{{\mathcal{B}}}{\times}{\mathcal{A}}\simeq{\mathcal{A}}\times\operatorname{Fib}({\mathcal{A}}\to{\mathcal{B}})$.
@@
There exists a monoidal functor
@@
8.8.11.
The constructions in Sect. 8.8.6 have immediate counterparts for the global Hecke stack
@@
Recall that the space $\operatorname{Ge}_{\pi_{1}({\check{G}})}(X)_{x}$ is canonically the discrete set $\pi_{1}({\check{G}})(-1)$ (see (8.36)), which is in bijection with the set of characters of $Z_{G}$. Denote the element corresponding to our gerbe by $\chi$.
8.9.2.
We claim now that the proof of the spectral decomposition theorem in [Ga4, Sect. 11.1] [^22] The proof in loc.cit. relies on two ingredients that were not written down at the time, but are available now. One is what is stated as [Ga4, Theorem 10.3.4], for which the proof has been supplied in [GLC2, Sects. 15-16]. Another is [Ga4, Proposition 11.1.3], for which the proof has been supplies in [FH] (in fact, as was observed by J. Faergeman, this second ingredient is not even necessary). applies in the current context. Recall that in loc. cit. the proof was based on considering the localization functor
@@
The key point in the proof of [Ga4, Theorem 4.5.2] is the fact (going back to [BD] and proved in [GLC2, Sects. 15-16]) that the functor
@@
Now, by the same principle as in [Ga4, Theorem 10.3.4] (see [GLC2, Sects. 15-16] for a full argument), we obtain that the functor $\operatorname{Loc}_{G_{\operatorname{ad}},\underline{x},\chi}$ factors as